# Activated LoRA: Fine-tuned LLMs for Intrinsics

### v2

@@ 3.2 Training and Huggingface Implementation
Our implementation of aLoRA is at https://github.ibm.com/Kristjan-H-Greenewald/activated-lora/tree/main. It seamlessly supports standard Huggingface training libraries Wolf et al. (2019) as well as inference/generation methods (with or wit hout using available base model KV cache) if needed (e.g. for testing or proofs of concept when the efficiencies of vLLM Zhu et al. (2023) are not needed). Later in this section, we will also show inference experiments using our modification of the more efficient (SOTA) inference package vLLM to support aLoRA.
@@
Training data is specified as a set of (possibly multiturn) inputs and single-turn completions, typically with the base model’s chat template applied. An invocation token sequence for the aLoRA model is optionally specified as described in the previous section. This invocation sequence is appended to the input sequence, and the model is finetuned to produce the output given the input.
To train the aLoRA adapter, a base model is specified, and we permit low-rank adapters to be applied to any (or all) query, key, and value blocks in the attention layers. [^1] We do not currently support adapting any other blocks. In training, the aLoRA is aware of which tokens occur before the invocation sequence, and does not adapt the weights for those tokens (as in (4)).
Generation with aLoRA is inherently simple due to the structure (4). When generating with the cache, the model simply uses the concatenation of the input base model KV cache with its own cache.
3.3 Model Capacity and Increased Rank
@@
In our experiments, we universally observed that aLoRA adapters required higher rank than LoRA adapters to get good performance. We offer the following intuition for this. LoRA adapters are free to adapt the weights for the keys and values for tokens prior to activation, so they are able to “compress” information needed for generation into the low-rank signal captured by the adaptation. This signal can then be “picked up” by the adapted query weights for generated tokens. On the other hand, aLoRA adapters are not able to do this compression, only being able to access base model KVs with adapted query weights. Hence, if the query adaptation is too low rank, it is not able span enough of the base model keys to capture all needed information. Figure 3 illustrates this. In our experiments, rank of 32 seems to be sufficient in most cases, which is still vastly smaller than the size of the base model. Whenever comparing to LoRA models, we chose a LoRA rank that achieved top performance for LoRA, rather than attempting to match the ranks between LoRA and aLoRA.
Note that the above provides the key motivation for starting the adapted weights at the intrinsic instruction tokens, rather than waiting to activate the weights only at generation time—this choice boosts model capacity.
(a) LoRA is able to modify the (K)Vs of the context to send (new) messages forward to the generation queries (Q).
(b) aLoRA is not able to do this (for the context), it can only look back at it with queries (Q).
Figure 3: ALoRA model capacity losses due to not being able to modify the keys and values of input tokens prior to the invocation sequence to send “messages” forward to the generation step. We will see empirically that any performance losses can be eliminated by increasing the rank of the adapter.
3.4 Comparison to LoRA at inference time
@@
To prove this point experimentally, we modified vLLM Zhu et al. (2023) to be able to do inference on activated LoRAs, trained four regular and activated LoRAs, and created a synthetic benchmark where the prompt length and the number of evaluators is varied. In Figure 5 we see the results of the experiment for the setting where the answer is comprised of 256 tokens and each evaluation is given by 16 tokens. The plots depict the total time processing the input prompts divided by the number of such prompts for IBM’s Granite 3.2 8 billion parameter instruct model (Granite 3.2 8b Instruct).
Note that the performance of activated LoRA changes very little as one adds more evaluators (more aLoRA adapted models in parallel), but the performance of regular LoRAs drops significantly as more adapters are added. This trend is consistent across all prompt lengths, showing very significant savings when using activated LoRA in this regime. We also emphasize that the use of aLoRA as an evaluator is for illustrative purposes; aLoRAs can also be used to generate steps in a reasoning process or a final answer in the case of more complex agent applications.
(a)promptanswereval1eval2prefill (base)gen (base)prefill (LoRA1)prefill (LoRA1) gen (LoRA1)prefill (LoRA2)prefill (LoRA2) gen (aLoRA2)(b)promptanswereval1eval2prefill (base)gen (base)gen (aLoRA1)gen (aLoRA2)
Figure 4: Computation and memory pattern of (a) LoRA vs. (b) aLoRA used as evaluators of an answer given by a base model. Narrow rectangles denote tokens and wide rectangles denote the KV cache. Significant GPU compute and memory savings can be realized because aLoRA can reuse the base model cache.
Figure 5: Comparison of aLoRA and LoRA as one varies the number of adapters, for a simple agent pattern. The experiment exemplifies that aLoRA scales much more efficiently than LoRA for this task.
4 Training adapters
In this section, we trained aLoRA adapters for IBM’s Granite 3.2 8b Instruct model, and compared their performance to performance achieved on trained LoRA models. As the LoRA models cannot re-use the base model cache, deploying them is not efficient, but serves as an accuracy benchmark to ensure aLoRA does not lose too much performance.
For aLoRA, all attention weights (keys, queries, values) were adapted in all layers, using rank 32 adapters. For both LoRA and aLoRA, the learning rate and number of epochs were tuned to achieve the best validation performance.
@@ 4 Training adapters
Note that since all the below models use the same base model (IBM’s Granite 3.2 8b Instruct), they can be swapped in and out as needed in the same flow. The reader can envision the possibilities of what can be built with these intrinsics, e.g. for RAG. See Danilevsky et al. (2025) for an exploration of composing these intrinsics (using LoRAs).
Comparing LoRA and aLoRA
To test the aLoRA framework, we consider several specific intrinsics tasks. Most of these intrinsics tasks were proposed or discussed in Danilevsky et al. (2025) , see that paper for deeper motivation for each intrinsic and extensive experimental results comparing standard LoRA implementations to other baselines. In this work, we therefore focus entirely on the comparison between LoRA and aLoRA performance, aiming to show that aLoRA models do not see meaningful degradation vs. LoRA models, while obtaining significant runtime benefits as discussed above.
See the appendices for additional details for the tasks and intrinsics.
4.1 Uncertainty Quantification
This intrinsic is designed to provide a Certainty score for model responses to user questions. The model will respond with a number from 0 to 9, corresponding to 5%, 15%, 25%,…95% confidence respectively. Training data for these confidence scores are obtained by applying the UQ calibration method of Shen et al. (2024a) to a large, diverse set of benchmark question answering datasets (see appendix) and quantizing the predicted confidences. Further details on the training process are in the appendix.
Certainty score interpretation
The returned percentages are calibrated in the following sense: given a set of answers assigned a certainty score of X%, approximately Xof these answers should be correct. Here “approximately" can be quantified via the expected calibration error, or ECE. Essentially what happens is teaching the adapter model what the base model knows and doesn’t know. This inherently requires generalization to questions of wildly varying difficulty (some of which may be trick questions!) and to settings not in training. Intuitively, it does this by extrapolating based on related questions it has been evaluated on in training - this is an inherently inexact process and leads to some hedging.
Usage
Note that the model is evaluating responses from its base model - in other words, it cannot be applied to generations from other models. The aLoRA architecture is thus particularly well-suited to this use case. In practice, the goal is to provide a highly efficient uncertainty score without having to resort to expensive larger judge models.
Metrics
The Uncertainty Quantification intrinsic returns an ordinal score, so we compute the mean absolute error between the predicted integer and the target integer in the training data.
Performance Results
Results are shown in Figure 6. Note that performance is largely unchanged using aLoRA instead of standard LoRA.
LoRA aLoRA
Uncertainty Quantification 0.50 MAE 0.49 MAE
@@ 4.1 Uncertainty Quantification
Figure 6: Comparison between LoRA and aLoRA test error for the Uncertainty Quantification intrinsic. Note that aLoRA does not lose meaningful performance.
4.2 Answerability Determination
This intrinsic is designed to assess whether a user’s final query in a multi-turn conversation can be answered given the retrieved documents. The model takes as input the full conversation history and a set of documents, and classifies the query as answerable or unanswerable based solely on the information contained in the set of input documents. This is useful in RAG settings. For instance, this intrinsic can help decide whether to proceed with generation or abstain with an “I don’t know” response.
The input to the model is a list of conversational turns and a list of documents converted to a string using apply_chat_template function. These turns can alternate between the user and assistant roles. The last turn is from the user. The list of documents is a dictionary with text field, which contains the text of the corresponding document. To prompt the aLoRA adapter to determine answerability, a special answerability role is used to trigger this capability of the model. The role includes the keyword "answerability": <|start_of_role|>answerability<|end_of_role|> When prompted with the above input, the model generates the answerable or unanswerable output. See Danilevsky et al. (2025) for more details.
Training Details
The aLoRA and LoRA adapters were fine-tuned using PEFT under the following regime: rank = 32, learning rate = 5e-6, number of epochs = 25, with early stopping based on validation set, and 90/10 split between training and validation.
Evaluation: Answerability Classification Performance
We evaluated aLoRA model against LoRA and baselines on binary answerability classification using two separate benchmarks:
  • Single-turn Setting (SQUADRun Benchmark Rajpurkar et al. (2018)): In this setting, the user query and the supporting documents are provided. Our model was evaluated against standard baselines to measure its ability to determine whether a standalone question is answerable based on the document set. Table 1 shows the classification results.
    Model Unans. Precision Unans. Recall Unans. F1 Ans. Precision Ans. Recall Ans. F1 Weighted F1
    Granite 3.2-8b LoRA 84.2 68.0 75.2 73.1 87.2 79.5 77.4
    Granite 3.2-8b aLoRA 83 81.1 82 81.4 83.3 82.4 82.2
@@ 4.2 Answerability Determination
    Table 1: Comparison of classification performance across models on SQUADRUN Dev set. Metrics are broken down by class (Answerable vs. Unanswerable) and include precision, recall, and F1 score.
  • Multi-turn Setting (MT-RAG Benchmark Katsis et al. (2025)): In this setting, the model is given the full multi-turn conversation history along with the supporting documents. This benchmark evaluates the model’s ability to assess answerability when the final user query can also depend on prior turns for context. Table 2 shows the results.
Model Unans. Precision Unans. Recall Unans. F1 Ans. Precision Ans. Recall Ans. F1 Weighted F1
Granite 3.2-8b LoRA 85.4 89.3 87.3 87.0 82.4 84.6 86.1
Granite 3.2-8b aLoRA 85.8 89.1 87.4 86.8 83 84.9 86.2
Table 2: Comparison of classification performance across models on MT-RAG Benchmark. Metrics are broken down by class (Answerable vs. Unanswerable) and include precision, recall, and F1 score.
Overall, the aLoRA model does not lose performance relative to the LoRA model, in fact, aLoRA achieves the best classification performance on SQUADRun. See Danilevsky et al. (2025) for additional baselines.
4.3 Query Rewrite
@@
This intrinsic is generally applicable for multi-turn conversational use cases, and its role is to perform rewrites of user queries for better performance for the downstream tasks. It is especially useful in RAG settings, see the metrics and evaluation results below. The query rewrite task is as follows: given a multi-turn conversation between a user and an AI assistant, de-contextualize the last user utterance (query) by rewriting it (whenever necessary) into an equivalent version that is standalone and can be understood by itself.
The rewrite is typically an expansion that in-lines, into the query, any implicit references that are made to entities, concepts, or even parts of the conversation that occur in the previous turns (either by the user or the AI assistant). Such expansion can include co-reference resolution (i.e., replacement of pronouns with the actual entities), handling of ellipsis, which is the common linguistic phenomenon where parts of a sentence or phrase are omitted by the user, but can be understood from the context (i.e., for whom, of what, with respect to something discussed above, etc.). The rewritten query can be sent to downstream tasks (e.g., to a retriever in a RAG setting) as a better replacement for the original user query, and without the need for (a potentially very long) context.
The input to the model is a list of conversational turns converted to a string using apply_chat_template function. These turns can alternate between the user and assistant roles, and the last turn is assumed to be from the user. To prompt the aLoRA adapter to rewrite the last user turn, a special rewrite role is used to trigger the rewrite capability of the model. The role includes the keyword "rewrite" followed by a short description of the query rewrite task.
Note: Even though one main application for query rewrite is in RAG settings, this intrinsic can be used to rewrite user questions for other conversational use cases (e.g., to access a database, or other APIs, or tools). As such, the adapter does not need any RAG documents (that may be present in the context, in a RAG setting) and uses only the dialog turns with what is being said between the user and assistant.
Training
Both the aLoRA and LoRA adapters were fine-tuned under the following regime: rank = 32, number of epochs = 25, with early stopping based on validation set, and 90/10 split between training and validation.
4.3.1 Query Rewrite: Evaluation on the Retrieval Task
We evaluate Recall@k on the MT-RAG benchmark Katsis et al. (2025), using either LoRA or aLoRA models to rewrite the query for the retriever. All retrieved passages are obtained using the Elser retriever with the same settings as in the above paper. We evaluate on three different testsets: a) full MT-RAG dataset (842 data points with last user turns); b) the non-standalone subset of MT-RAG dataset, which is a subset of 260 (out of 842) last user turns that were annotated by humans as non-standalone (i.e., they are dependent on the prior context); c) the standalone subset of MT-RAG dataset, which is the complementary subset, with all the last user turns that were annotated by humans as standalone.
@@ 4.3.1 Query Rewrite: Evaluation on the Retrieval Task
Retrieval recall evaluation (Recall@k) with different query rewrite strategies, evaluated on full, non-standalone and standalone subsets of MT-RAG dataset are shown in Tables 3, 4, and 5 respectively.
Rewrite Strategy Recall@5 Recall@10 Recall@20
aLoRA 0.54 0.66 0.74
LoRA 0.56 0.68 0.76
Table 3: Query rewrite strategies on the retrieval task of full MT-RAG dataset
Rewrite Strategy Recall@5 Recall@10 Recall@20
aLoRA 0.42 0.54 0.64
LoRA 0.44 0.57 0.66
Table 4: Query rewrite strategies on the retrieval task of non-standalone subset of MT-RAG
Rewrite Strategy Recall@5 Recall@10 Recall@20
aLoRA 0.63 0.75 0.82
LoRA 0.63 0.75 0.83
Table 5: Query rewrite strategies on the retrieval task of standalone subset of MT-RAG
@@
Note that throughout, performance numbers for aLoRA and LoRA are within a point or two. See CITERAG for additional comparisons showing that these approaches outperform benchmarks and are very close to the performance with gold rewrites.
4.3.2 Query Rewrite: Evaluation on Answer Generation
We evaluate answer generation quality, with top-k passages retrieved under the various query rewrite strategies for the retriever. We choose here $k=20$, but similar trends take place for other values of $k$. We used Granite-3.2-8b instruct as the answer generator, and RAGAS Faithfulness (RAGAS-F) and RAD-Bench score as metrics for answer quality. We use the same three testsets as above.
The answer quality evaluation using RAGAS-F and RAD-Bench on full, non-standalone and standalone subsets of MT-RAG dataset are shown in Tables 6, 7, and 8 respectively.
Rewrite Strategy RAGAS-F RAD-Bench
aLoRA 0.81 0.69
LoRA 0.81 0.70
Table 6: Impact of query rewrite strategies on answer generation on full MT-RAG dataset
Rewrite Strategy RAGAS-F RAD-Bench
aLoRA 0.77 0.69
LoRA 0.79 0.69
Table 7: Impact of query rewrite strategies on answer generation on non-standalone subset of MT-RAG
Rewrite Strategy RAGAS-F RAD-Bench
aLoRA 0.83 0.70
LoRA 0.83 0.71
Table 8: Impact of query rewrite strategies on answer generation on standalone subset of MT-RAG
As with Recall, observe that LoRA and aLoRA performance is very close, indicating that moving to aLoRA for inference reasons does not require giving up any performance.
4.4 Jailbreak Detection
This intrinsic is designed for detecting jailbreak risk within user prompts. Prompts with jailbreak risk vary across a wide range of attack styles - from direct instructions, to encoding-style, social-hacking based attacks and even ones that exploit special token or context overload (Rawat et al., 2024). In our experiments we focused on training intrinsics for detecting social hacking style of adversarial prompts. The intrinsic is trained to return a binary label - “Y” indicating jailbreak risk present and “N” indicating no risk, when invoked. We evaluate jailbreak intrinsic across a mixture of samples with jailbreak risk (Shen et al., 2024b; Lin et al., 2023; Wan et al., 2024) and benign samples (Conover et al., 2023). As with the other intrinsics, we see very little performance difference between LoRA and aLoRA.
Jailbreak Risk Detector Acc TPR FPR
@@ 4.4 Jailbreak Detection
LoRA 0.944 0.898 0.009
Table 9: Performance of jailbreak risk detectors across an aggregated mixture of prompts with jailbreak risk and benign samples.

### v4

@@ 2.1 Training
Our implementation of aLoRA is in the supplementary material. It seamlessly supports standard Huggingface training libraries Wolf et al. (2019) as well as inference/generation methods (with or without using available base model KV cache) if needed (e.g. for testing or proofs of concept when the efficiencies of vLLM (Zhu et al., 2023) are not needed). Later in this section, we will also show inference experiments using our modification of the more efficient (SOTA) inference package vLLM to support aLoRA.
@@
For aLoRA SFT, training data is specified as a set of (possibly multiturn) inputs and completions, typically with the base model’s chat template applied. An invocation token sequence for the aLoRA model is optionally specified as described in the previous section. This invocation sequence is appended to the input sequence, and the model is finetuned to produce the output given the input. To train the aLoRA adapter, a base model is specified, and following standard LoRA practice, we apply low-rank adapters to any (or all) of the query, key, and value blocks in the attention layers. [^4] Adapting MLP layers is possible but less common in LoRA SFT. In training, the aLoRA is aware of which tokens occur before the invocation sequence, and does not adapt the weights for those tokens (as in (6)).
@@
Adapter Model Capacity and Increased Rank: Empirically, we observe that aLoRA adapters sometimes require higher rank than LoRA adapters to get good performance (no experiments in this paper use rank higher than $r=32$). We offer the following intuition for this. LoRA adapters are free to adapt the weights for the keys and values for tokens prior to activation, so they are able to “compress” information needed for generation into the low-rank signal captured by the adaptation. This signal can then be more easily “picked up” by the adapted query weights for generated tokens. See Appendix D for further exploration. This also provides the key motivation for starting the adapted weights at the intrinsic instruction tokens, rather than waiting to activate the weights only at generation time—this choice boosts model capacity.
We now compare aLoRA to standard LoRA both in terms of inference compute costs as well as generation quality and accuracy. Recall that our goal is to achieve significant gains in inference cost, while preserving (not losing) accuracy relative to what is achievable with LoRA.
3 Inference timing results on vLLM
Figure 3: Comparison of aLoRA and LoRA when used as evaluators in a simple agentic pattern. Top left: Multiplicative speedup of an aLoRA evaluator vs LoRA, showing up to 35x improvement depending on base model and prompt length. Top right: Multiplicative speedup for the end-to-end pipeline including the base model generation (256 tokens) and 1 or 5 parallel eval (adapter) generations (16 tokens each). Despite the large fixed cost of the base model call, end-to-end aLoRA speedups are still significant, highlighting LoRA inefficiency. Bottom row: Log-log plots for the wall clock evaluation, showing that even for small models, the delay for LoRA becomes meaningful in absolute terms as the prompt and number of evaluations in the agentic pipeline scale.
@@ 3 Inference timing results on vLLM
To prove this experimentally, we modified vLLM (Zhu et al., 2023) to be able to perform inference on aLoRAs with base model prefix cache reuse, and compared LoRA and aLoRA inference on 7 small ($\leq 14$B) LLMs in the setting where the initial base model prompt length is varied, the base model generated answer is comprised of 256 tokens, and each adapter evaluation is given by 16 tokens. Results are in Figure 3, showing timing for both 1 evaluator and 5 evaluators. [^5] While 5 evaluators may initially seem a lot, this is a stand-in for more complex pipelines where the input, retrieved documents, and generation are all being checked, as in Danilevsky et al. (2025).
The speedup factor of aLoRA increases as prompt length increases, adapter generation time improving over 2 to 7$\times$ even for 250-token prompts, and over 20$\times$ for most models on the longest prompts. Despite most of the compute advantages coming from the prefill savings (time to 1st token of the adapter), we still see highly meaningful speedups in the end-to-end pipeline which also includes the large fixed cost of generating 256 tokens with the base model. Finally, in the last row, despite the adapters only generating 16 tokens, we see that the LoRA delays are noticeable in wall clock time. Note these wall clock delays will be further magnified for more complex agentic or reasoning pipelines where many rounds of input and generation happen before the overall pipeline completes.
4 Finetuning Accuracy Results
Having demonstrated significant inference time speedups for aLoRA, it remains to ascertain whether we lose generation accuracy or quality. We first train LoRA and aLoRA adapters on 4 LLMs on a set of benchmark SFT tasks, and then consider a set of more challenging “intrinsics”-type tasks for which well-engineered, very recent LoRA adapters already exist against which we can compare.
4.1 Benchmark SFT tasks
Task Llama 3.2 1B Llama 3.2 3B Llama 3.1 8B Mistral 7B
LoRA aLoRA LoRA aLoRA LoRA aLoRA LoRA aLoRA
Bengali Hate Speech Classification 79.30% 81.94% 86.34% 89.43% 70.04% 85.02% 72.25% 85.46%
WIQA: Effect Classification 68.92% 71.38% 76.15% 76.00% 74.92% 78.00% 61.08% 79.08%
MMLU Conceptual Physics MCQA 33.33% 38.89% 72.20% 66.67% 55.56% 55.56% 55.56% 55.56%
MMLU College Computer Science MCQA 66.67% 58.33% 66.67% 75.00% 66.67% 58.33% 75.00% 75.00%
SocialIQA Question Generation 86.00% 88.77% 89.85% 90.15% 52.00% 88.92% 97.23% 90.92%
Hindi Sentence Perturbation 69.60% 74.69% 98.30% 63.89% 86.11% 35.19% 99.23% 96.30%
SuperGLUE Question Generation 98.42% 95.79% 95.26% 96.84% 98.95% 92.11% 99.47% 92.11%
Figure 4: LoRA vs. aLoRA accuracy (%) on each task across base models after hyperparameter grid search guided by the validation set. While individual task performance is noisy due to the size of the datasets etc., there is no consistent accuracy loss from using aLoRA over LoRA.
In Brüel-Gabrielsson et al. (2024), a collection of 1000 instruction SFT tasks were curated and LoRA adapters were trained for each. These tasks were drawn from from the Super-Natural Instructions Wang et al. (2022a) benchmark collection of datasets, which in turn drew from sources such as MMLU Hendrycks et al. (2021) etc.
From this list, we selected 7 tasks at random, roughly split between multiple-choice and freeform outputs. We excluded from consideration any tasks with (a) too small datasets, (b) overly open-ended instructions, or (c) difficulty such that trained adapters did no better than random chance. The selected tasks and datasets are detailed in Appendix F. For each task, we trained both LoRA and aLoRA adapters for 4 instruct-tuned models of various sizes. To ensure a fair comparison, for each task-model pair, using the validation set performance we performed grid search over 4 rank values and 5 learning rate values [^6] Note that this experiment represents 1120 separate adapter training runs. and followed typical LoRA SFT best practices for all parameters. The invocation sequence was set to equal the model’s chat template generation prompt. See Appendix F for further details.
Results are shown in Figure 4. While there is decent variability due to the generally small size of the datasets and non-exhaustiveness of the hyperparameter search, overall it can be seen that neither aLoRA or LoRA has a consistent accuracy advantage. The median performance difference is 0.0%, and the mean performance difference is 0.6% in favor of LoRA (3.06% in favor of activated LoRA when restricted to MC tasks). Using a 2 sided t-test, the p-value for rejecting the null hypothesis of equal means is 0.8, failing to reject at level 0.05. This supports our thesis that aLoRA can achieve comparable accuracy to LoRA while significantly outperforming in inference costs.
4.2 Intrinsics Tasks
In this subsection, we consider real-world tasks that fit the framework of intrinsics, specifically, settings where the task requires a potentially long input that either (a) already has base model KV cache available from either being input to or generated by the base model, or (b) does not have (complete) base model KV cache available, but base model KV cache created through an aLoRA call will be used in a subsequent base model call. We draw from those proposed in Danilevsky et al. (2025) for a RAG pipeline, though these intrinsics are not limited to RAG. See Danilevsky et al. (2025) for deeper motivation for each intrinsic and extensive experimental results comparing their LoRA adapters to strong baselines.
@@ 4.2 Intrinsics Tasks
In Danilevsky et al. (2025), LoRA adapters were trained for the intrinsics using the Granite 3.2 8b Instruct model. For direct comparison to their released adapters, we train corresponding aLoRA adapters (using the same datasets and chat-template-based formatting) and compare their performance in terms of generation quality and/or accuracy. [^7] As they released LoRA adapters for only one base model, we limit ourselves to their choice of base model. Note that since all the below models use the same base model, they can be swapped in and out as needed in the same flow. The reader can envision the possibilities of what can be built with these intrinsics, e.g. for RAG. See the App. G for additional details for this section.
Certainty Score LoRA aLoRA
MAE 0.50 0.49
@@
Figure 5: Test error for the Uncertainty Quantification intrinsic. Note that aLoRA does not lose meaningful performance.
Uncertainty Quantification: This intrinsic provides a Certainty score for model responses to user questions. The model will respond with a number from 0 to 9, corresponding to $5\%,15\%,25\%,\dotsc,95\%$ confidence respectively. Training data for these confidence scores are obtained by applying the UQ calibration method of Shen et al. (2024a) to a large, diverse set of benchmark question answering datasets, quantizing the predicted confidences and using these predicted confidences as SFT targets for the adapter. Following the chat template, the invocation sequence is 4 tokens: <|start_of_role|>certainty<|end_of_role|>. Note that the model is evaluating responses from its base model - in other words, it cannot be applied to generations from other models. The aLoRA architecture is thus particularly well-suited to this use case. In practice, the goal is to provide a highly efficient uncertainty score without having to resort to expensive larger judge models.
The Uncertainty Quantification intrinsic returns an ordinal score, so we compute the mean absolute error between the predicted integer and the target integer in the SFT data. Results are shown in Figure 5. Note that performance is largely unchanged using aLoRA instead of standard LoRA.
Dataset Adapter Unans. Ans. Weighted F1
P R F1 P R F1
SQUADRUN Dev LoRA 84.2 68.0 75.2 73.1 87.2 79.5 77.4
aLoRA 83.0 81.1 82.0 81.4 83.3 82.4 82.2
MT-RAG Benchmark LoRA 85.4 89.3 87.3 87.0 82.4 84.6 86.1
aLoRA 85.8 89.1 87.4 86.8 83.0 84.9 86.2
@@
Figure 6: Comparison of classification performance of LoRA vs. aLoRA across the SQUADRUN Dev set and MT-RAG benchmark. Metrics are broken down by class (Unanswerable vs. Answerable) and include precision (P), recall (R), F1 score, and weighted F1.
Answerability Determination: This intrinsic assesses whether a user’s final query in a multi-turn conversation can be answered given the retrieved documents. In RAG or other settings, this decides whether to proceed with generation or abstain with an “I don’t know” response. The invocation sequence is set to <|start_of_role|>answerability<|end_of_role|>. We tested on binary answerability classification on the single-turn SQUADRun Benchmark Rajpurkar et al. (2018) with the user query and the supporting documents, and the multi-turn MT-RAG Benchmark Katsis et al. (2025) using full multi-turn conversation history along with the supporting documents. Figure 6 shows the results. Overall, the aLoRA model does not lose performance relative to the LoRA model.
Query Rewrite:
@@
This intrinsic is generally applicable for multi-turn conversational use cases, and its role is to perform rewrites of user queries for better performance for the downstream tasks. It is especially useful in RAG settings, see the metrics and evaluation results below. The query rewrite task is as follows: given a multi-turn conversation between a user and an AI assistant, de-contextualize the last user utterance (query) by rewriting it (whenever necessary) into an equivalent version that is standalone and can be understood by itself. The rewritten query can be sent to downstream tasks (e.g., to a retriever in a RAG setting) as a better replacement for the original user query, and without the need for (a potentially very long) context.
Full MT-RAG Non-standalone Standalone
Strategy R@5 R@10 R@20 R@5 R@10 R@20 R@5 R@10 R@20
aLoRA 0.54 0.66 0.74 0.42 0.54 0.64 0.63 0.75 0.82
LoRA 0.56 0.68 0.76 0.44 0.57 0.66 0.63 0.75 0.83
(a) Retrieval (Recall@5, @10, and @20)
Full MT-RAG Non-standalone Standalone
Strategy RAGAS-F RAD-Bench RAGAS-F RAD-Bench RAGAS-F RAD-Bench
aLoRA 0.81 0.69 0.77 0.69 0.83 0.70
LoRA 0.81 0.70 0.79 0.69 0.83 0.71
(b) Answer generation quality (RAGAS-F, RAD-Bench)
Figure 7: Impact of query rewrite strategies on both retrieval and generation tasks across subsets of MT-RAG.
@@
Retrieval recall evaluation (Recall@k) with different query rewrite strategies, evaluated on full, non-standalone and standalone subsets of MT-RAG dataset Katsis et al. (2025) are shown in Figure 1(a). All retrieved passages are obtained using the Elser retriever with the same settings as in the above paper. We evaluate on three different subsets of MT-RAG detailed in the appendix.
@@
Answer quality evaluation using RAGAS Faithfulness (RAGAS-F) and RAD-Bench on full, non-standalone and standalone subsets of MT-RAG dataset are shown in Table 1(b) (see appendix for details). Note that throughout, performance numbers for aLoRA and LoRA are within a point or two. See Danilevsky et al. (2025) for additional comparisons showing that these approaches outperform benchmarks and are very close to the performance with gold rewrites.
Acc TPR FPR
@@
LoRA 0.943 0.898 0.011
Figure 8: Performance of jailbreak risk detectors.
Jailbreak Detection: This intrinsic is designed for detecting jailbreak risk within user prompts. Prompts with jailbreak risk vary across a wide range of attack styles - from direct instructions, to encoding-style, social-hacking based attacks and even ones that exploit special token or context overload (Rawat et al., 2024). In our experiments we focused on training intrinsics for detecting social hacking style of adversarial prompts. As with prior intrinsics, the aLoRA/LoRA detectors are trained in the same conditions with a rank of 32. The intrinsic is trained to return a binary label - “Y” indicating jailbreak risk present and “N” indicating no risk.
After training, we evaluate the jailbreak intrinsic on new out-of-distribution datasets not used in training to robustly assess generalizability Zizzo et al. (2025). The out-of-distribution datasets comprise of 3,282 samples containing a mixture of samples with jailbreak risk (Shen et al., 2024b; Lin et al., 2023; Wan et al., 2024) and benign samples (Conover et al., 2023). As with the other intrinsics, we see very little performance difference between LoRA and aLoRA.